Please help me!
The length of a rectangle is y + 5, and the width of the rectangle is y - 3. What is the area of the rectangle? Formula: A = lw

Answer by fractalier(6550)

Area A = LW = (y+5)(y-3) = y^2 + 5y - 3y - 15 = y^2 + 2y - 15
done by the FOIL Method...